# HeurekaBench: A Benchmarking Framework for AI Co-scientist

### v1

@@ A.4 Additional details on question filtering
Our question generation pipeline converts validated insights into questions. However, since we utilize LLMs for this task, it remains essential to perform additional filtering to develop a more robust benchmark. For the first automatic filtering stage, we rely on closed-source LLMs. We use both GPT-4o and Claude-4-Sonnet to answer all OEQs and MCQs. For MCQs, we discard any questions that both GPT-4o and Claude-4-Sonnet correctly answered. For OEQs, we use G-Eval (liu2023geval) to assign a score between 1 and 5, and eliminate questions that received scores above 3.0 for both LLMs. This aims to reduce the risk of LLM-based agents answering from pre-training knowledge. The second manual filtering stage first checks for potential hallucinations and duplication by cross-referencing with the insights representation. Next, for insights containing multiple findings but validation only for a subset (e.g., if the experimental techniques suggest immunofluorescence staining was used by authors to validate some part of the insight, we mark that part as not validated), we remove questions derived from the non-validated components.
@@ Appendix B Details on Evaluation in HeurekaBench framework
One of our key contributions is the adaptation of G-Eval (liu2023geval) to evaluate data-driven agent responses against ground-truth conclusions derived from scientific findings, while simultaneously penalizing responses that rely predominantly on pre-training knowledge of the LLM rather than exploring the provided dataset. Moreover, the generated response may itself contain novel findings, an essential aspect of the AI co-scientist, when it is tasked with discovering patterns from experimental data. To support this, we design a grading rubric grounded in atomic facts extracted from free-form text. We define atomic facts as minimal, verifiable units, such as conditions, trends, and conclusions, that can be systematically compared across texts. Concretely, we instruct the GPT-4o grader to first decompose responses into atomic facts, and then assess whether each ground-truth fact is fully present, partially present, or absent in the agent’s answer. The score increases with the extent and number of correctly captured atomic facts, provided that no statement contradicts the ground truth. The full evaluation prompt is included below.
@@ C.1 Selected publications for Insight generation
Lazarescu2025AdipocyteAtlas Nature Genetics 4 4
Yu2025NeuroblastomaAtlas Nature Genetics 8 8
Maatz2025CardiacResponses Nature Cardiovascular Research 1 1
Wang2025NeocortexDynamics Nature 0 1
Zhang2024LimbAtlas Nature 2 3
Gu2024ImmuneMicroniches Nature 2 2
Kedlian2024MuscleAgingAtlas Nature Aging 11 10
Zwick2024EpithelialZonation Nature Cell Biology 2 2
Isola2024AgingOvary Nature Aging 6 5
Rexach2024DementiaGenomics Cell 3 3
Hoo2024PlacentaResponse Cell Systems 5 7
Escoubas2024Microglia Cell 2 1
Li2024UterineNK Cell Stem Cell 4 3
@@ C.3 Categorization of open-ended questions (OEQs)
(vi) cell-cell communication analysis is specific to the final publication (Li2024UterineNK), which deals with cell-cell communication between uNK and trophoblast cells (e.g., “How do uterine natural killer (uNK) cell-derived cytokines influence the expression of cytokine receptors in extravillous trophoblasts (EVTs)?”).
@@
Figure 4: Question distribution per task category. There are six different categories, with the distribution across both versions of the benchmark following a similar distribution. The questions related to cell-cell communication analysis are obtained only from Li2024UterineNK, which is not part of the sc-HeurekaBench-Lite.
@@ C.4 Invalidated insights
As a result of this validation step, some insights are invalidated. The primary reasons include: (i) insufficient information about the dataset used to derive the insight. In such cases, running the workflow on an alternative dataset produces results that differ from those reported in the paper. For example, the workflow specifies bulk RNA-seq data, while only scRNA-seq data is available from CellxGene (czi2025cz). (ii) inconsistencies between the workflow’s requirements and the available dataset. For instance, a workflow may reference sub-cluster information that the dataset does not provide. (iii) overly generic insights that cannot lead to meaningful questions (e.g., The study presents a comprehensive human skeletal muscle aging atlas, providing a valuable resource for studying muscle aging across species.). While we mark the above insights as invalid in our benchmark, we expect that these could potentially be validated if the correct set of input information is provided.
@@ C.5 Comparative Analysis with BaisBench
In this section, we perform a comparative analysis with BaisBench (luo2025benchmarking). BaisBench generates only multiple-choice questions (MCQs) from scientific publications using a single LLM, which (i) do not undergo any validation, and (ii) do not use the corresponding experimental data or code repositories for grounding the questions. In contrast, our proposed HeurekaBench utilizes scientific publications along with their corresponding code repositories to ensure that each question-answer pair (i) depends on manually validated insights through code execution, and (ii) is grounded in corresponding experimental data.
Furthermore, we perform quantitative analysis to compare the quality of the generated questions. We use GPT-5 (openai2025gpt5) to answer MCQs from both datasets. In this setup, the LLM is given access only to questions, not to any experimental data. Hence, it tests the data-driven nature of the questions, i.e., whether they can be answered correctly based on knowledge recall. GPT-5 answered 53.37% of questions accurately on BaisBench, compared to 34.69% on sc-HeurekaBench, indicating that our framework yields questions that are more difficult to answer with recall alone.
@@ D.1 Scraping publications and expert insights from FlyBase
To curate pairs of expert insights and publications from FlyBase (ozturk2024flybase), we focus on the “Other Comments” section of gene-level report pages. We select the following gene identifiers: FBgn0000490, FBgn0001077, FBgn0001139, FBgn0001168, FBgn0001180, FBgn0003448, FBgn0003731, FBgn0003996, FBgn0004644, FBgn0004647. We collect five random expert insights with open-access publications per identifier, resulting in a total of 50 pairs. Within each insight, we append the abbreviated gene names with their full names from the gene report pages for additional context to LLM-Judge.
@@ D.2 GPT-4o judge for matching our and expert insights
To validate our InsightExtractor, we rely on G-Eval (liu2023geval) to compare if our generated insights from the publication resemble the expert-annotated findings from the corresponding dataset. We instruct the G-Eval to label relatedness as strongly related, weakly related, or unrelated. The entire prompt is available below.
@@ D.3 Manual Analysis of AI Co-scientist Answers
Figure 5: Score distribution per task category. Both versions of the benchmark exhibit a similar trend in score distributions across three evaluated LLMs, with the closed-source model outperforming the open-source model in all categories. Cell–cell communication category is sourced exclusively from Li2024UterineNK, which is not included in sc-HeurekaBench-Lite.
@@ D.4 Inter-rater agreement statistics
As detailed in Section 4.3.1, we perform inter-rater agreement studies between three closed-source LLM judges, GPT-4o, Claude-4.5-Sonnet (anthropic2025claude45), and Gemini-2.5-Pro (comanici2025gemini). We computed Spearman’s rank correlation and the unweighted Cohen’s kappa ($\kappa$) with a quadratic penalty between the correctness scores assigned by GPT-4o and those assigned by other judges. We report these two statistics for three planner LLMs (Qwen3-235B-thinking, GPT-OSS-120B, Claude-4-Sonnet) within the Biomni agent in Table 6.

### v2

@@ A.4 Additional details on question filtering
Our question generation pipeline converts validated insights into questions. However, since we utilize LLMs for this task, it remains essential to perform additional filtering to develop a more robust benchmark. For the first automatic filtering stage, we rely on closed-source LLMs. We use both GPT-4o and Claude-4-Sonnet to answer all OEQs and MCQs. For MCQs, we discard any questions that both GPT-4o and Claude-4-Sonnet correctly answered. For OEQs, we use G-Eval (Liu et al., 2023) to assign a score between 1 and 5, and eliminate questions that received scores above 3.0 for both LLMs. This aims to reduce the risk of LLM-based agents answering from pre-training knowledge. The second manual filtering stage first checks for potential hallucinations and duplication by cross-referencing with the insights representation. Next, for insights containing multiple findings but validation only for a subset (e.g., if the experimental techniques suggest immunofluorescence staining was used by authors to validate some part of the insight, we mark that part as not validated), we remove questions derived from the non-validated components.
@@ Appendix B Details on Evaluation in HeurekaBench framework
One of our key contributions is the adaptation of G-Eval (Liu et al., 2023) to evaluate data-driven agent responses against ground-truth conclusions derived from scientific findings, while simultaneously penalizing responses that rely predominantly on pre-training knowledge of the LLM rather than exploring the provided dataset. Moreover, the generated response may itself contain novel findings, an essential aspect of the AI co-scientist, when it is tasked with discovering patterns from experimental data. To support this, we design a grading rubric grounded in atomic facts extracted from free-form text. We define atomic facts as minimal, verifiable units, such as conditions, trends, and conclusions, that can be systematically compared across texts. Concretely, we instruct the GPT-4o grader to first decompose responses into atomic facts, and then assess whether each ground-truth fact is fully present, partially present, or absent in the agent’s answer. The score increases with the extent and number of correctly captured atomic facts, provided that no statement contradicts the ground truth. The full evaluation prompt is included below.
@@ C.1 Selected publications for Insight generation
Lazarescu et al. (2025) Nature Genetics 4 4
Yu et al. (2025) Nature Genetics 8 8
Maatz et al. (2025) Nature Cardiovascular Research 1 1
Wang et al. (2025) Nature 0 1
Zhang et al. (2024) Nature 2 3
Gu et al. (2024b) Nature 2 2
Kedlian et al. (2024) Nature Aging 11 10
Zwick et al. (2024) Nature Cell Biology 2 2
Isola et al. (2024) Nature Aging 6 5
Rexach et al. (2024) Cell 3 3
Hoo et al. (2024) Cell Systems 5 7
Escoubas et al. (2024) Cell 2 1
Li et al. (2024) Cell Stem Cell 4 3
@@ C.3 Categorization of open-ended questions (OEQs)
(vi) cell-cell communication analysis is specific to the final publication (Li et al., 2024), which deals with cell-cell communication between uNK and trophoblast cells (e.g., “How do uterine natural killer (uNK) cell-derived cytokines influence the expression of cytokine receptors in extravillous trophoblasts (EVTs)?”).
@@
Figure 4: Question distribution per task category. There are six different categories, with the distribution across both versions of the benchmark following a similar distribution. The questions related to cell-cell communication analysis are obtained only from Li et al. (2024), which is not part of the sc-HeurekaBench-Lite.
@@ C.4 Invalidated insights
As a result of this validation step, some insights are invalidated. The primary reasons include: (i) insufficient information about the dataset used to derive the insight. In such cases, running the workflow on an alternative dataset produces results that differ from those reported in the paper. For example, the workflow specifies bulk RNA-seq data, while only scRNA-seq data is available from CellxGene (CZI Cell Science Program et al., 2025). (ii) inconsistencies between the workflow’s requirements and the available dataset. For instance, a workflow may reference sub-cluster information that the dataset does not provide. (iii) overly generic insights that cannot lead to meaningful questions (e.g., The study presents a comprehensive human skeletal muscle aging atlas, providing a valuable resource for studying muscle aging across species.). While we mark the above insights as invalid in our benchmark, we expect that these could potentially be validated if the correct set of input information is provided.
@@ C.5 Comparative Analysis with BaisBench
In this section, we perform a comparative analysis with BaisBench (Luo et al., 2025). BaisBench generates only multiple-choice questions (MCQs) from scientific publications using a single LLM, which (i) do not undergo any validation, and (ii) do not use the corresponding experimental data or code repositories for grounding the questions. In contrast, our proposed HeurekaBench utilizes scientific publications along with their corresponding code repositories to ensure that each question-answer pair (i) depends on manually validated insights through code execution, and (ii) is grounded in corresponding experimental data.
Furthermore, we perform quantitative analysis to compare the quality of the generated questions. We use GPT-5 (OpenAI, 2025) to answer MCQs from both datasets. In this setup, the LLM is given access only to questions, not to any experimental data. Hence, it tests the data-driven nature of the questions, i.e., whether they can be answered correctly based on knowledge recall. GPT-5 answered 53.37% of questions accurately on BaisBench, compared to 34.69% on sc-HeurekaBench, indicating that our framework yields questions that are more difficult to answer with recall alone.
@@ D.1 Scraping publications and expert insights from FlyBase
To curate pairs of expert insights and publications from FlyBase (Öztürk-Çolak et al., 2024), we focus on the “Other Comments” section of gene-level report pages. We select the following gene identifiers: FBgn0000490, FBgn0001077, FBgn0001139, FBgn0001168, FBgn0001180, FBgn0003448, FBgn0003731, FBgn0003996, FBgn0004644, FBgn0004647. We collect five random expert insights with open-access publications per identifier, resulting in a total of 50 pairs. Within each insight, we append the abbreviated gene names with their full names from the gene report pages for additional context to LLM-Judge.
@@ D.2 GPT-4o judge for matching our and expert insights
To validate our InsightExtractor, we rely on G-Eval (Liu et al., 2023) to compare if our generated insights from the publication resemble the expert-annotated findings from the corresponding dataset. We instruct the G-Eval to label relatedness as strongly related, weakly related, or unrelated. The entire prompt is available below.
@@ D.3 Manual Analysis of AI Co-scientist Answers
Figure 5: Score distribution per task category. Both versions of the benchmark exhibit a similar trend in score distributions across three evaluated LLMs, with the closed-source model outperforming the open-source model in all categories. Cell–cell communication category is sourced exclusively from Li et al. (2024), which is not included in sc-HeurekaBench-Lite.
@@ D.4 Inter-rater agreement statistics
As detailed in Section 4.3.1, we perform inter-rater agreement studies between three closed-source LLM judges, GPT-4o, Claude-4.5-Sonnet (Anthropic, 2025a), and Gemini-2.5-Pro (Comanici et al., 2025). We computed Spearman’s rank correlation and the unweighted Cohen’s kappa ($\kappa$) with a quadratic penalty between the correctness scores assigned by GPT-4o and those assigned by other judges. We report these two statistics for three planner LLMs (Qwen3-235B-thinking, GPT-OSS-120B, Claude-4-Sonnet) within the Biomni agent in Table 6.